### v5

Abstract
We bridge quasi-experimental and structural approaches to develop a robust method for retrospective merger evaluation. First, we show that the standard difference-in-differences estimation equation can be interpreted as a “reduced form” of our structural equations, formally bridging these methodologies. Demand and cost parameters enable the construction of a robust measure of the merger price effect that remains identified even when the difference-in-differences approach is not, and can be decomposed into efficiency gains versus conduct changes among the remaining firms. Second, we enhance the structural approach by weighting each market’s contribution to the moment conditions using market- and time-specific synthetic weights when the number of treated markets is limited. Applying this methodology to three major airline mergers, we find that modest efficiency gains are entirely offset by increased coordination, particularly in later mergers. The synthetic GMM refinement sharpens these findings, uncovering anti-competitive effects that standard approaches miss.
JEL: L40, L41, D43
Keywords: Merger Analysis, Airline Competition, Antitrust Policy, Synthetic GMM.
1 Introduction
Merger evaluation is an important area of research in industrial organization and antitrust economics. The challenge in merger evaluation lies in distinguishing between mergers that generate efficiency and those that primarily increase market concentration, enabling firms to exercise market power (Farrell and Shapiro, 1990) and facilitate coordination (Miller and Weinberg, 2017; Porter, 2020). Two methodological approaches dominate this field: quasi-experimental methods, particularly difference-in-differences (DiD) that compare outcomes between treated and control markets (Kim and Singal, 1993; Ashenfelter and Hosken, 2010; Carlton et al., 2019), and structural approaches that estimate demand and cost parameters to simulate merger outcomes (Nevo, 2000). [^1] For a comprehensive overview see Whinston (2006); Kwoka, Jr. (2014) and Asker and Nocke (2021).
Quasi-experimental methods aim to offer transparent identification and deliver easily interpretable merger effects. Structural methods emphasize the underlying economic mechanisms that determine the unilateral effects of a merger. Each approach addresses different aspects of merger evaluation. The “credibility revolution” (Angrist and Pischke, 2010; Nevo and Whinston, 2010) highlights the potential complementarity between these approaches. However, they remain largely disconnected in merger evaluation, often yielding conflicting conclusions about the same mergers (Peters, 2006; Björnerstedt and Verboven, 2016).
In this paper, we develop a two-way methodological bridge between these two approaches for robust retrospective merger evaluation. For the first bridge, we show that the DiD specification is the “reduced form” of demand and supply equations, clarifying when DiD may fail to identify merger effects, while the structural approach remains reliable. For the second bridge, we propose modifying GMM by weighting each observation’s contribution to the moment conditions using synthetic DiD weights–synthetic GMM. These weights select better control markets, improving the precision and credibility of structural estimation.
We apply our method to evaluate three major airline mergers: Delta-Northwest (2008), United-Continental (2010), and American-US Airways (2013). Airline industry has been widely studied in academia (Borenstein, 1990; Kim and Singal, 1993; Peters, 2006; Olley and Town, 2018; Carlton et al., 2019; Das, 2019; Li et al., 2022; Eizenberg and Zvuluni, 2024; Olsen et al., 2024; Remer and Orchinik, 2024; Bruegge et al., 2024) and legal proceedings (U.S. District Court, MA, 2023).
The paper is organized into three parts. First, we start with a quasi-experimental analysis that reveals how methodological choices fundamentally determine merger evaluation conclusions. Standard DiD with fixed effects suggests that all three mergers reduced prices by 4-8% and expanded capacity by 16-29%. However, these pro-competitive effects vanish when using the first-difference method. This striking sensitivity of the DiD approach motivates our methodological innovations. [^2] Although our contexts differ, de Chaisemartin and D’Haultfœuille (2020) also suggest that the difference between two-way fixed effects and first-difference estimators can inform about the data-generating process.
Second, we develop our first bridge by showing that the DiD is the reduced form of structural demand and supply equations. This connection reveals why DiD may not identify merger effects because, unlike standard policy interventions, mergers involve strategic interactions among firms that correlate with unobserved (to researchers) factors that affect demand, induce competitor responses, and may even alter competitive conduct—each potentially invalidating DiD identification, consistent with Marx et al. (2024). Our structural framework addresses these limitations by explicitly modeling these economic forces and relying on instrumental variables for identification.
We model firm behavior using a conduct parameter, following Bresnahan (1982), and allow it to change in response to a merger. If the conduct parameter increases, then the merger has a coordination effect. With appropriate instruments, such as cost shifters (potential competitor presence, operational characteristics) and demand shifters (network effects, market size), we can identify merger-induced efficiency and the change in conduct but not levels. Using our structural model, we can also construct a DiD-like but robust measure of the aggregate price effect of a merger, which has an intuitive economic interpretation: it equals efficiency gains adjusted by the cost pass-through rate.
Our estimates reveal a concerning temporal progression. The DL-NW merger exhibits decreased coordination (0.141-0.141), but this change is statistically insignificant. However, subsequent mergers increase coordination, with AA-US demonstrating significant competitive deterioration as its conduct increases by $0.107$ (5% level). This escalating pattern validates the DOJ’s concerns about coordinated effects that led it to block the AA-US merger. These conduct changes largely offset modest efficiency gains, yielding neutral price effects. However, several of the estimates are imprecise because there are limited treated markets–a common challenge in merger evaluation, where only a handful of markets or units are typically treated.
Third, we develop our second bridge to utilize the DiD approach in addressing such imprecision in structural estimates. In particular, we propose using market- and time-specific weights from synthetic DiD (Arkhangelsky et al., 2021) to weight each observation’s contribution to the moment conditions. These weights help make treated and control markets comparable (Abadie and Gardeazabal, 2003), which should improve the precision of the “synthetic GMM” estimates. [^3] While our synthetic GMM approach is motivated specifically by merger analysis with limited treatment markets, the general principle of reweighting observations to address sampling concerns has precedent. For instance, Manski and Lerman (1977); Manski and McFadden (1981); Imbens (1992) proposed weighting likelihood and moment contributions to correct for choice-based sampling bias in discrete choice models. Indeed, we find that after weighting, estimates have better precision. Moreover, they suggest anti-competitive impacts that standard approaches miss: UA-CO shows significant coordination increases (0.049, 10% level) with efficiency losses of \$13.5 per \$100 ticket, suggesting that the price increases price by \$12.9, representing a 5.7% increase relative to the sample mean of \$228 for treated markets before meger, while AA-US exhibits substantial coordination effects (0.100, 5% level) with no discernable price effects.
Our two-way bridge between quasi-experimental and structural approaches speaks to longstanding tensions in empirical economics between credible identification and economic insight (Leamer, 1983; Bresnahan, 1989). As merger enforcement continues to evolve and face scrutiny (Kwoka, Jr., 2014; Scott Morton, 2019; Rose and Shapiro, 2022; Kaplow, 2024), developing empirical methods that combine the strengths of multiple approaches while addressing their limitations becomes increasingly important for informed antitrust policy.
2 Data
We use data from three sources. First, we use the US Department of Transportation’s Origin and Destination Survey (DB1B), a quarterly 10% sample of all tickets sold, which contains information on ticket prices and passenger numbers for each carrier and route. Second, the Bureau of Transportation Statistics’ T100 (Form 41) database is used to obtain the number of seats (capacity) allocated by each airline across all routes on a quarterly basis. Third, we also utilize the Official Airline Guide (OAG) database for all nonstop flight schedules.
We focus on three major mergers in the airline industry. Delta and Northwest Airlines (DL-NW) announced their merger plan on April 14, 2008; the shareholders approved it on September 26, 2008, and the US Department of Justice (DOJ) approved it on October 29, 2008. United and Continental (UA-CO) announced their merger plan on May 2, 2010, gaining shareholder approval on September 17, 2010. The merger was completed on October 1, 2010, forming United Continental Holdings; however, the full operational integration was not completed until mid-2012. The merger between American Airlines and US Airways (AA-US) was announced on February 14, 2013. This merger faced initial opposition from the DOJ, which filed a lawsuit on August 13, 2013, to block the merger. However, a settlement was reached on November 12, and the merger was completed on December 9, 2013.
We adopt the framework established by Carlton et al. (2019) to define our analytical parameters, including time, control, and treatment groups. Specifically, for each merger, the pre- and post-merger periods are defined, respectively, as the eight quarters preceding and following the merger approval. Specifically, for DL-NW, these are Q4 2006-Q3 2008 (pre) and Q1 2009-Q4 2010 (post); for UA-CO, these are Q3 2008-Q2 2010 (pre) and Q4 2010-Q3 2012 (post); and for AA-US, these are Q4 2011-Q3 2013 (pre) and Q1 2014-Q4 2015 (post).
Next, we discuss the selection of “treated” and “control” markets, an approach that the US court system accepts when evaluating merger cases. For each merger, we classify a city-pair market as a “treated” market if both merging firms provided nonstop service. [^4] A carrier is considered ‘nonstop’ in a city route if it conducts 10 nonstop operations (five round-trips).
To minimize any indirect effects of the merger from these selected markets, we exclude markets where either of the two merging parties had a substantial connecting presence. A carrier is deemed to have a substantial connecting presence in a market if it handles at least 10% of total passenger traffic through connections. Specifically, we remove all “connecting overlap” markets, where, prior to the merger, each party served at least 10% of passengers and jointly served at least 40% of passengers. We refer to the remaining markets as the treated markets.
Likewise, we classify a market as a “control” market if the merger does not affect its market structure. In other words, in such a market, only one carrier can be the merging party, so if it is a duopoly or a triopoly before the merger, it remains the same after the merger. Thus, the control markets have the same number of nonstop carriers as the overlap routes in the pre-merger period. As with the treated market, we also restrict our attention to control markets where, whenever applicable, the merging firm does not have a substantial connecting presence.
Table 1: Summary Statistics
Pre-Merger Post-Merger
Mean Median SD Mean Median SD
Average Price
   Treated 228.25 229.52 52.24 226.33 231.52 48.32
   Control 185.65 179.05 52.70 195.75 189.65 51.93
   Treated and Control 186.22 179.63 52.92 196.16 190.19 51.99
   Top 50 MSAs 190.37 183.93 51.99 199.99 193.51 51.35
Passengers
   Treated 37,546.17 29,755 25,643.97 39,261.17 30,420 27,690.30
   Control 34,330.75 25,760 32,748.04 33,884.38 25,130 32,655.25
   Treated and Control 34,374.05 25,850 32,664.22 33,956.57 25,190 32,598.85
   Top 50 MSAs 41,753.65 32,110 34,851.46 41,760.55 32,620 34,710.27
Offered Seats
   Treated 132,391.60 155,774.50 70,204.90 148,652.90 170,601.50 81,357.05
   Control 103,468.40 76,187.50 88,269.59 100,320.80 74,366.00 86,191.09
   Treated and Control 103,858.20 76,558.50 88,112.42 100,970.10 74,749.50 86,305.49
   Top 50 MSAs 120,719.80 94,867.00 95,021.14 118,584.60 93,900.00 92,958.21
  • Note: This table shows the mean, median, and standard deviation for our main variables, broken out by different market types (all markets, treated and control markets, and top 50 MSA markets) and merger status.
In Table 1, we present summary statistics revealing distinct patterns across market segments before and after the merger. Despite maintaining higher average fares, treated markets showed a slight decrease from \$228.25 to \$226.33 post-merger. Notably, control markets experienced a price increase from \$185.65 to \$195.75. The average one-way fare in the combined sample of treated and control markets and in the sample of markets between the top 50 MSAs increased, respectively, from \$186.22 to \$196.16 and from \$190.37 to \$199.99.
Passenger volumes exhibited mixed trends: while the combined sample of treated and control markets saw a slight decline from 34,374 to 33,956 passengers, treated markets experienced growth from 37,546 to 39,261 passengers, representing approximately a 4.6% increase. Finally, the seat capacity in treated markets grew substantially from 132,391 to 148,652 seats (a 12.3% increase), while the overall market average decreased from 103,858 to 100,970 seats.
3 Quasi-Experimental Approach
We use an event-study approach to assess the impact of three airline mergers (Delta-Northwest, United-Continental, and American-US Airways) on fares, passenger volume, and seat availability using a difference-in-differences (DiD) strategy. The DiD strategy compares the difference in the average outcome before and after a merger for treated markets with the difference for control markets to get the average effect of a merger on the treated markets.
3.1 Empirical Specification
We begin by estimating the following model:
$Y_{mt}=\beta_{0}+\beta^{\texttt{DiD}}\times I_{mt}+X_{mt}^{\top}\gamma+\delta_{m}+\delta_{t}+\varepsilon_{mt},$ (1)
where $I_{mt}:=\text{Treated}_{m}\times\text{Post-Merger}_{mt}$ is the treatment indicator variable with $\text{Treated}_{m}$ equals one if $m$ is a treated market and zero otherwise, and $\text{Post-Merger}_{mt}$ equals one if $t$ is after a merger and zero otherwise, and $Y_{mt}$ is the outcome variable (logarithm of average nominal fares, passenger volume, or seat availability) for market $m$ in year-quarter $t$. The specification includes time-varying controls $X_{mt}$ (e.g., share of nonstop passengers), market fixed effects, $\delta_{m}$, and time fixed effects, $\delta_{t}$.
To identify the effect of a merger, we need three assumptions. First, the no anticipation assumption ensures that future mergers do not affect pre-merger outcomes. Second, the Stable Unit Treatment Value Assumption (SUTVA) requires that potential outcomes for market $m$ depend only on $m$’s treatment status, i.e., if it is affected by a merger, ruling out spillover across markets. Third, the parallel trends assumption requires that the treated and the control markets would have followed the same trend in the absence of the merger.
Under our specification in Equation (1), potential outcomes, as a function of $I_{mt}$, are:
$Y_{mt}(0)$ $=$ $\beta_{0}+X_{mt}^{\top}\gamma+\delta_{m}+\delta_{t}+\varepsilon_{mt};$
$Y_{mt}(1)$ $=$ $\beta_{0}+\beta^{\texttt{DiD}}+X_{mt}^{\top}\gamma+\delta_{m}+\delta_{t}+\varepsilon_{mt}.$
Then, the parallel trends assumption implies that counterfactual trends are identical across treated and control markets, i.e., $\mathbb{E}[Y_{m,\text{post}}(0)-Y_{m,\text{pre}}(0)|\text{Treated}_{m}=1]=\mathbb{E}[Y_{m,\text{post}}(0)-Y_{m,\text{pre}}(0)|\text{Treated}_{m}=0].$ Using the potential outcome framework and following standard derivations, this assumption reduces to the condition on unobservables
$\mathbb{E}[\varepsilon_{m,\text{post}}-\varepsilon_{m,\text{pre}}|\text{Treated}_{m}=1]=\mathbb{E}[\varepsilon_{m,\text{post}}-\varepsilon_{m,\text{pre}}|\text{Treated}_{m}=0],$ (2)
provided that observable controls, $X_{mt}$, evolve similarly across the two groups of markets. This condition requires that the expected change in unobservable factors be identical for both groups while allowing for different levels of unobservables between groups or periods.
Under these assumptions, the regression coefficient $\beta^{\text{DiD}}$ in (1) identifies the average treatment effect of the merger on the treated markets as:
$\beta^{\text{DiD}}$ $=$ $\Big(\mathbb{E}[Y_{m,\text{post}}|\text{Treated}_{m}=1]-\mathbb{E}[Y_{m,\text{post}}|\text{Treated}_{m}=0]\Big)$
$-\Big(\mathbb{E}[Y_{m,\text{pre}}|\text{Treated}_{m}=1]-\mathbb{E}[Y_{m,\text{pre}}|\text{Treated}_{m}=0]\Big).$
The first difference captures the post-merger gap between treated and control markets, and the second difference captures the pre-merger gap. By subtracting the pre-treatment difference, DiD removes time-invariant differences between groups, identifying the merger effect.
3.2 DiD Estimates of Merger Effects
We begin by estimating Equation (1) using fixed effects, the standard approach in the literature. Table 2 presents the DiD estimates of $\beta^{\text{DiD}}$. The results suggest substantial pro-competitive effects across all mergers. Average prices decreased by 4.02% for DL-NW, 3.54% for UA-CO, and 7.51% for AA-US. [^5] The DiD estimate can be interpreted as a percentage change in the outcome variable, where the percentage change for $\beta^{\texttt{DiD}}$ is given by $(\exp(\beta^{\texttt{DiD}})-1)\times 100$ (Halvorsen and Palmquist, 1980). Passenger volume increased by 2.84%, 5.87%, and 9.20% respectively, though the DL-NW effect is imprecisely estimated. Most notably, the number of offered seats expanded dramatically: 24.12% for DL-NW, 29.43% for UA-CO, and 16.53% for AA-US.
Table 2: Estimated Effects of Mergers
Outcomes DL-NW UA-CO AA-US
Log (Price) $041^{***} 0.0$$36^{***} 0.0$$078^{***}$
$(0.015)$ $(0.008)$ $(0.025)$
Log (Passengers) $0.028$ $057^{***} 0.0$$088^{***}$
$(0.018)$ $(0.010)$ $(0.030)$
Log (Offered Seats) $0.216^{***}$ $0.258^{***}$ $0.153^{***}$
$(0.047)$ $(0.009)$ $(0.031)$
  • Note: This table reports DiD estimates from Equation (1) for airfare, passenger volume and offered seats using fixed effects. The analysis examines the three major airline mergers separately: Delta-Northwest (DL-NW), United-Continental (UA-CO), and American-US Airways (AA-US). $^{***}:p<0.01$.
These findings suggest all three mergers improved consumer welfare through lower prices and expanded service. Our results align with Carlton et al. (2019), who used similar DiD fixed-effects specifications and concluded that these mergers generated consumer benefits.
3.3 Identification of DiD with Fixed Effects
While the DiD framework provides a useful approach for evaluating mergers, its identification is tenuous in merger contexts. To understand these challenges systematically, we can conceptualize the DiD “error” term in Equation (1) $\varepsilon_{mt}$ as reflecting underlying economic forces that drive market outcomes. As it will become clearer later when we study the structural approach in Section 4, this error term can be decomposed into demand-side and supply-side components, making the connection between DiD identification failures and economic fundamentals transparent.
Endogenous Selection. The first concern arises from the strategic nature of merger decisions. The parallel trends assumption essentially requires that, conditional on observables ($X_{mt}$) and fixed effects ($\delta_{m}$, $\delta_{t}$), the selection of treated markets is “as good as random” with respect to future outcome trends. However, airlines evaluate potential merger partners based on network characteristics and anticipated market trajectories, creating two types of selection. Selection on demand occurs when merger decisions correlate with demand trends—airlines might merge in declining markets or, conversely, in growing markets to capitalize on expansion opportunities. Selection on supply occurs when merger decisions correlate with cost trajectories—airlines might target markets where joint operations would reduce costs or where standalone operations face rising expenses. Either form of selection invalidates the parallel trends assumption.
Differential Market Characteristics. Even with exogenous merger motivations—driven by non-market factors such as financial considerations or regulatory constraints—the parallel trends assumption can still fail. If merging airlines systematically serve markets with different underlying demand and cost trajectories, then treated and control markets will diverge even in the absence of strategic selection.
Strict Exogeneity. An additional concern arises from the temporal structure of merger analysis. The fixed effects DiD estimator requires strict exogeneity: $\mathbb{E}[\varepsilon_{mt}\times I_{ms}]=0$ for all $m,t,s$. This assumption is particularly fraught when markets exhibit differential time trends. Consider the case where $\varepsilon_{mt}=\alpha_{m}+\kappa_{m}\cdot t+\nu_{mt}$, with $\nu_{mt}$ representing idiosyncratic noise and $\kappa_{m}$ capturing market-specific linear trends. If treated and control markets have different trend parameters, i.e., $\mathbb{E}[\kappa_{m}|\mathrm{Treated}_{m}=1]\neq\mathbb{E}[\kappa_{m}|\mathrm{Treated}_{m}=0]$, then bias grows linearly with time.
Specifically, when using a four-quarter pre- and post-merger window, the average bias across post-merger quarters equals $2.5(\kappa_{1}-\kappa_{0})$, where $\kappa_{1}$ and $\kappa_{0}$ represent average trend parameters for treated and control markets, respectively. Therefore, even a slight deviation from the pre-existing trend can result in substantial errors.
3.4 DiD with Market Trends
We conducted pre-trend analysis and found systematic pre-trends across all outcome variables—airfares, passenger volumes, and offered seats—consistent with the identification concerns raised above. [^6] For brevity, we do not report the full pre-trend analysis here, as this represents standard event study methodology. Complete pre-trend analysis results are available upon request. These findings suggest including market-specific linear time trends in our DiD specification to control for systematic differences in market trajectories that would otherwise invalidate the parallel trends assumption.
To this end, we model the error term in (1) as $\varepsilon_{mt}=\alpha_{m}+\kappa_{m}\times t+\nu_{mt}$, where $\kappa_{m}\times t$ captures market-specific linear trends, and estimate the following model:
$Y_{mt}=\beta_{0}+\beta^{\texttt{DiD}}\times I_{mt}+X_{mt}^{\top}\gamma+\delta_{m}+\delta_{t}+\kappa_{m}\times t+\nu_{mt}.$ (3)
We implement two methods. Method 1 uses the entire sample period to estimate market-specific trends. Method 2 uses a two-step procedure that first estimates trends using only pre-merger data, and then extrapolates these trends to estimate the merger effects on detrended outcomes. For each method, we consider two variants: differential trends between treated and control groups (Panel A) and market-specific trends (Panel B). Table 3 shows the results.
Table 3: Estimated Effects of Mergers, with Time Trends
Method 1 Method 2
DL-NW UA-CO AA-US DL-NW UA-CO AA-US
Panel A: Treatment-Control Specific Trends
Log (Price) 0.022-0.022 $074^{***} 0.0$320.032 $041^{***} 0.0$$52^{***} 0.0$$063^{**}$
$(0.023)$ $(0.020)$ $(0.057)$ $(0.015)$ $(0.008)$ $(0.025)$
Log (Passengers) $0.162^{***}$ $105^{***} 0.0$800.080 $105^{***} 0.0$10-0.010 $0.148^{***}$
$(0.027)$ $(0.012)$ $(0.063)$ $(0.018)$ $(0.010)$ $(0.030)$
Log (Offered Seats) $156^{***} 0.0$$063^{***}$ $0.123^{**}$ $0.354^{***}$ $0.160^{***}$ $0.193^{***}$
$(0.037)$ $(0.014)$ $(0.050)$ $(0.047)$ $(0.009)$ $(0.031)$
Panel B: Market-Specific Trends
Log (Price) 0.023-0.023 $073^{***} 0.0$310.031 $0.045^{**}$ $^{***} 0.1210.121^{**}$
$(0.024)$ $(0.020)$ $(0.057)$ $(0.019)$ $(0.046)$ $(0.061)$
Log (Passengers) $0.161^{***}$ $104^{***} 0.0$820.082 $0.101^{*}$ $0.120^{**}$ $0.243^{***}$
$(0.028)$ $(0.012)$ $(0.063)$ $(0.052)$ $(0.048)$ $(0.106)$
Log (Offered Seats) $151^{***} 0.0$$062^{***}$ $0.124^{**}$ $0.306^{***}$ $0.124^{***}$ $0.287^{***}$
$(0.038)$ $(0.014)$ $(0.048)$ $(0.045)$ $(0.035)$ $(0.103)$
  • Note: This table reports DiD estimates of merger effects from Equation (3) controlling for time trends. Panel A controls for differential trends between treatment and control groups; Panel B allows for market-specific trends. Method 1 estimates trends jointly with merger effects using the full sample. Method 2 estimates trends using pre-merger data only, then applies DiD to detrended outcomes. Standard errors clustered at the market level are in parentheses. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
Adding time trends substantially alters the merger effects. The negative price effects from Table 2 either reverse sign or become statistically insignificant. This reversal suggests that the estimated price reductions in Table 2 reflected pre-existing upward price trends in control markets rather than merger-induced efficiencies.
The previously positive passenger effects either disappear or reverse. Individual merger results show mixed patterns: DL-NW consistently shows significant passenger decreases, while AA-US shows increases in some specifications. However, the positive effect on seat capacity remains robust across all specifications, though magnitudes vary considerably. In comparison, baseline estimates indicated a 16.53% to 29.43% capacity increase, and trend-adjusted estimates (Method 1) range from 6.39% to 16.88%. Method 2 consistently produces larger capacity effects than Method 1.
The persistence of capacity expansion, coupled with the reversal in price and passenger effects, suggests changes in airline conduct rather than efficiency gains. The increase in offered seats, combined with insignificant or negative passenger effects, implies substantial reductions in load factors. This pattern may indicate strategic capacity deployment to maintain market presence rather than demand-driven expansion. The positive price effects in many trend-adjusted specifications are consistent with this interpretation.
While trend-adjusted results provide valuable insights, they rely on linear time trend specifications that may be inadequate when markets follow nonlinear trajectories. Hub markets may experience accelerating growth due to network effects, leisure destinations might exhibit cyclical patterns, or mature markets could display diminishing growth rates. Controlling for linear trends may not capture differential market evolution, and the parallel trends assumption may still fail. This limitation motivates consideration of first differencing, which accommodates arbitrary nonlinear trends under weaker identifying assumptions.
3.5 DiD Estimates using First Differences
First differencing offers several advantages over the previous approaches. We find that outcome variables are non-stationary in levels but become stationary (Im et al., 2003) after first-differencing, addressing potential spurious regression concerns. More importantly, first differences require only that period-to-period changes in trends be similar across treated and control markets, accommodating arbitrary nonlinear trend differences between groups.
However, using first differences to estimate merger effects introduces a new challenge. Standard DiD estimates of (1) using the first-difference method capture only immediate post-merger effects, as the first-difference of the treatment indicator $\Delta I_{mt}$ equals one only in the period immediately following the merger. This timing limitation is particularly problematic for mergers, where the post-merger integration of operations and network optimization may not be immediate. Consequently, the realization of cost efficiencies may unfold gradually over multiple quarters or years following the merger. This pattern has been documented across several industries (Focarelli and Panetta, 2003; Gugler and Siebert, 2007; Ashenfelter and Hosken, 2010; Ashenfelter et al., 2013) and is likely applicable to airline mergers as well.
To address this limitation while maintaining the advantages of first differences, we allow for the gradual realization of merger effects over time. We estimate:
$Y_{mt}=\beta_{0}+\sum_{k>0}^{T}\beta^{\texttt{DiD}}_{k}\times I_{m,t,k}+X_{mt}^{\top}\gamma+\delta_{m}+\delta_{t}+\varepsilon_{mt},$
where $I_{m,t,k}=\text{Post-Merger}_{t,k}\times\text{Treated}_{m}$ and $\text{Post-Merger}_{t,k}$ equals one if by period $t$ it has been at least $k$ periods since the merger. This modelling approach allows each $\beta^{\texttt{DiD}}_{k}$ to vary over time, enabling us to trace how merger effects evolve as integration progresses.
Taking first differences, we estimate:
$\Delta Y_{mt}$ $=$ $\sum_{k=1}^{T}\beta^{\texttt{DiD}}_{k}\times\Delta I_{m,t,k}+\Delta X_{mt}^{\top}\gamma+\Delta\delta_{t}+\Delta\varepsilon_{mt},$ (4)
where
[…]
The parallel trends assumption for first differences requires that counterfactual trends be identical across treated and control markets: $\mathbb{E}[\Delta\varepsilon_{mt}|\text{Treated}_{m}=1]=\mathbb{E}[\Delta\varepsilon_{mt}|\text{Treated}_{m}=0]$. While fixed effects require that the change in trend levels be the same, first differences require that period-to-period changes in trends be the same, which is substantially weaker. This approach accommodates arbitrary nonlinear trends, including different trend levels, growth rates, or functional forms, as long as the period-to-period changes are similar across groups.
Additionally, first differencing relaxes the strict exogeneity requirement to sequential exogeneity, mitigating the bias amplification problem associated with multi-period analysis windows. Our approach thus combines the robustness of first differences to nonlinear trends with the economic realism of gradual merger implementation.
We extend our analysis to $T=8$ quarters (two years) post-merger to balance the tradeoff between allowing sufficient time for efficiency gains to materialize and avoiding contamination from other market developments. This horizon maintains comparability with the fixed effects estimates in Table 2.
Figure 1: Estimated Effects of Airline Mergers over Time (updated)
Note: Note: Figure displays first-difference DiD estimates of merger effects on log prices (red), log passenger volume (blue), and log seats (green) for three airline mergers. Each bar represents the estimated coefficient for the treatment effect $k$ quarters after merger completion defined in (4) and (5), with 95% confidence intervals shown as error bars. Robust standard errors are clustered at the city-route level.
Figure 1 presents the first-difference estimates of the merger effects with time-varying treatment effects, allowing us to track how merger impacts evolve as integration progresses. Each “Quarter k” captures the effect when a treated market transitions into the post-merger state in quarter $k$ following merger completion. This specification reveals how merger effects unfold over the course of twelve quarters following each merger.
We find little evidence of stable price reductions across any of the three mergers. The DL-NW and UA-CO mergers exhibit modest price impacts that remain largely neutral over the observation period, while the AA-US merger shows no clear downward trend. Thus any efficiency benefits from these mergers do not translate into meaningful price reductions.
Table 4: Estimated Merger Effects
DL-NW UA-CO AA-US
Price Pass. Seats Price Pass. Seats Price Pass. Seats
0.0097 -0.0170 0.1838*** -0.0168** -0.0146 0.1782*** 0.0053 0.1283** 0.1518***
[0.610] [0.475] [0.000] [0.040] [0.548] [0.000] [0.898] [0.029] [0.000]
  • Note: Table reports average merger effects on log outcomes estimated over eight post-merger quarters (detailed quarterly estimates shown in Figure 1). Each transformed value is computed as a weighted average of $(\exp(\beta_{k})-1)$, where $\beta_{k}$ is the estimated log-point effect in post-merger quarter $k$. This transformation maps log changes into percentage changes in original units. The weighting scheme gives more weight to periods closer to the merger date, with weights declining linearly over time. Standard errors are computed using the delta method applied to the variance-covariance matrix of the underlying quarterly estimates.
    $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
We assess this impression by examining average effects over different time horizons. Table 4 reveals that airline mergers generally fail to deliver the promised efficiency gains that would benefit consumers through lower prices during the critical two-year post-merger integration period. Across the first eight quarters, two of the three mergers (DL-NW and AA-US) exhibit neutral price effects, while only UA-CO shows a meaningful price reduction of 1.68%. The DL-NW merger demonstrates no significant price change (0.97% increase) over quarters 1–8, while the AA-US merger similarly shows neutral effects (0.53% increase).
The heterogeneous outcomes underscore that merger effects depend critically on market-specific factors and the integration strategies employed, challenging the efficiency rationale. Before examining the underlying mechanisms that drive these diverse outcomes, we summarize our findings thus far.
Our quasi-experimental analysis reveals that methodological choices fundamentally determine merger evaluation conclusions. Standard DiD estimates suggest uniformly pro-competitive effects, yet these findings either disappear entirely or attenuate and become mixed when accounting for pre-existing market trends or using the first-difference method.
Beyond these methodological concerns, quasi-experimental approaches face deeper conceptual limitations. DiD estimates a single average treatment effect, masking rich heterogeneity in merger impacts and providing no guidance about which mergers benefit consumers or under what conditions. We therefore turn to a structural approach to address these limitations and provide economic insights into merger mechanisms.
4 Structural Approach
This section develops a structural framework to evaluate mergers that captures key economic mechanisms affecting merger outcomes. With this approach, we can determine not just whether prices change, but also why. In particular, it allows us to decompose the price effect of a merger into efficiency, market power, and coordination effects.
Our first bridge demonstrates that the canonical DiD specification in Equation (1) emerges as a reduced-form of the equilibrium demand and supply conditions. This connection also clarifies why DiD identification may fail in merger settings and enables us to construct an aggregate measure of price effects analogous to the DiD coefficient, using demand and cost parameters, that is more robust than the DiD estimate. Our second bridge employs recent advances in synthetic control and synthetic DiD methods to improve control market selection, improving the precision of structural estimates.
4.1 Demand and Supply
To maintain consistency with our quasi-experimental approach, we model demand for air travel at the market level, treating air travel as a homogeneous product sold by $n$ symmetric firms. The demand depends on price and exogenous variables, and is given by $Q=D(P,Z;\boldsymbol{\alpha})+\varphi$, where $Q$ is the quantity demanded, $P$ is the market price, and $Z$ represents exogenous factors that shift demand but are unaffected by either price $P$ or the demand “shock” $\varphi$. Here, $\boldsymbol{\alpha}$ is the vector of parameters that determine the demand function.
On the supply side, we model marginal cost as $C=c(Q,W,I;\boldsymbol{\theta})+\vartheta$, where $Q$ is quantity supplied, $W$ is exogenous cost-shifters independent of $P$, cost shock $\vartheta$ and demand shock $\varphi$, and $I\in\{0,1\}$ is the merger indicator equaling one post-merger in treated market and zero otherwise, and $\boldsymbol{\theta}$ is the vector of cost parameters. Including merger $I$ in the cost function captures merger-induced efficiency, for instance, if $c(Q,W,1;\boldsymbol{\theta})\leq c(Q,W,0;\boldsymbol{\theta})$.
Suppressing the demand and cost parameters, firm $i$’s profit function is given by $\Pi_{i}=P(Q)q_{i}-c(q_{i})$ and firm $i$ chooses $q_{i}$ to maximize its profit, where $P(Q)$ is the inverse demand function. To capture competitive interactions, we follow Bresnahan (1982) and include a strategic conduct parameter $\lambda\in[0,1]$ that measures the intensity of competition. In other words, if $\{q_{i}\}_{i=1,\ldots,N}$ is equilibrium output with total output $Q=\sum_{i}q_{i}$ if for each firm $i$ its output $q_{i}=\arg\max_{\tilde{q}_{i}}\left\{P(\lambda(\tilde{q}_{i}-q_{i})+Q)\tilde{q}-c(\tilde{q}_{i})\right\}$, such that if firm $i$ deviates from $q_{i}$ by $(\tilde{q}_{i}-q_{i})$ other firms will respond by changing the total output by $\lambda\times(\tilde{q}_{i}-q_{i})$. Therefore, this conduct parameter $\lambda$ measures market competitiveness, with $\lambda=0$ representing perfect competition, $\lambda=1$ indicating monopoly, and intermediate values reflecting varying degrees of competitive intensity. Then, the marginal revenue function at the market level becomes:
$MR=P+\lambda\times\left(\frac{\partial P}{\partial Q}\times Q\right)=P+\lambda\times h(Q),$
where $h(Q)\leq 0$ represents the standard markup term from the inverse demand function.
Mergers can also affect competitive conduct. For instance, sustaining supracompetitive pricing may become easier when fewer firms remain post-merger, as monitoring becomes easier, which in turn facilitates tacit coordination. This aspect of the merger suggests conduct may shift toward monopolistic behavior following a merger. To capture this effect, we allow the conduct to change with the merger, such that $\lambda(I=0)=\lambda_{\text{pre}}\leq\lambda(I=1)=\lambda_{\text{post}}$.
Incorporating this merger-dependent conduct parameter into the pricing equation and equating marginal revenue with marginal cost yields the equilibrium pricing function:
$P=c(Q,W,I;\boldsymbol{\theta})-\lambda(I)\times h(Q,Y,\boldsymbol{\alpha})+\vartheta.$ (6)
Thus, mergers can affect equilibrium prices through cost efficiency and strategic conduct.
We assume that mergers affect only the supply-side relationships but not the demand. This assumption reflects the principle that preferences remain constant following mergers, even as consumers adjust to the resulting price changes. Although firms may strategically modify their product attributes after a merger, such as route networks, service quality, or operational frequencies, we assume that the unobservable determinants of these adjustments are statistically independent of the demand and cost shocks in pricing equations. [^7] This orthogonality condition is implicit in the DiD framework. When estimating demand and cost parameters, we use only the characteristics of competing non-merging firms, excluding those of merging firms, to minimize the effects of product repositioning that may occur after the merger.
4.2 Identification
To ensure comparability with the linear DiD specification, we impose additional structure on the demand and cost parameters. Specifically, we assume linearity:
$Q_{mt}$ $=$ $D(P_{mt},Z_{mt};\boldsymbol{\alpha})+\varphi_{mt}=\alpha_{0}+\alpha_{1}P_{mt}+\alpha_{2}Z_{mt}+\varphi_{mt};$
$C_{mt}$ $=$ $c(Q_{mt},W_{mt},I_{mt};\boldsymbol{\theta})=\theta_{0}+\theta_{1}Q_{mt}+\theta_{2}W_{mt}+\theta_{3}I_{mt}+\vartheta_{mt},$
where $\alpha_{1}<0$ ensures the demand decreases with prices.
These linear specifications, combined with the first-order optimality condition from Equation (6) and the relationship $h(Q)=Q_{mt}/\alpha_{1}$, yield a system of simultaneous equations governing equilibrium outcomes in each market $m$ during period $t$:
$Q_{mt}$ $=\alpha_{0}+\alpha_{1}P_{mt}+\alpha_{2}Z_{mt}+\varphi_{mt}$ (7)
$P_{mt}$ $=\theta_{0}+\tilde{\gamma}_{mt}Q_{mt}+\theta_{2}W_{mt}+\theta_{3}I_{mt}+\vartheta_{mt}$
where the (nuisance) parameter $\tilde{\gamma}_{mt}:=(\theta_{1}-\lambda(I_{mt})/\alpha_{1})$ captures the economies of scale $\theta_{1}$, strategic conduct $\lambda(I_{mt})$ and the demand slope $\alpha_{1}$. [^8] As Abbring et al. (2025) point out, our structural approach traces back to Wright (1928), who introduced structural models of demand and supply, established identification using instrumental variables, and developed methods for counterfactual analysis nearly a century ago. For the ease of presentation, using the definition of $I_{mt}$ from (1), let $\tilde{\gamma}_{mt}=\tilde{\gamma}_{1}$ if $I_{mt}=1$ and $\tilde{\gamma}_{mt}=\tilde{\gamma}_{0}$ otherwise, and with a slight abuse of notation, we continue to use $\tilde{\boldsymbol{\theta}}$ to represent the cost parameters $(\theta_{0},\tilde{\gamma}_{1},\tilde{\gamma}_{0},\theta_{2},\theta_{3})^{\top}$.
4.2.1 Demand and Supply Parameters
Our identification strategy relies on carefully constructed exclusion restrictions that can help “trace” demand and supply functions. Specifically, we require that cost-shifting variables $W_{mt}$ influence consumer demand only indirectly through their effect on market prices, while demand-shifting variables $Z_{mt}$ can affect prices only indirectly through their effect on demand. In particular, we assume the following moment restrictions:
$\text{Cov}(W_{mt},\varphi_{mt})$ $=0\quad\text{(Cost shifters orthogonal to demand shocks)},$ (8)
$\text{Cov}(Z_{mt},\vartheta_{mt})$ $=0\quad\text{(Demand shifters orthogonal to supply shocks)}.$
These conditions combined with instrument relevance requirements, i.e., $\mathbb{E}[P_{mt}W_{mt}]\neq 0$ and $\mathbb{E}[Q_{mt}Z_{mt}]\neq 0$, ensure that the instrumental variables $(Z,W)$ provide sufficient conditions to identify a subset of demand and cost parameters.
Lemma 1.
Under the orthogonality conditions in (8) and instrument relevance assumptions, the parameters ${\boldsymbol{\alpha}}:=(\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$ of the demand and $\tilde{\boldsymbol{\theta}}:=(\theta_{0},\tilde{\gamma}_{1},\tilde{\gamma}_{0},\theta_{2},\theta_{3})^{\top}$ of the supply relation defined in Equation (7) are identified.
The proof is in the Appendix B. While the demand and supply relations in (7) are written for one-dimensional instruments for expositional clarity, in our proof and when estimating, we use multiple IVs for each Equation. For example, for the airline industry, cost-shifters include potential competitor presence at origin and destination airports, legacy carrier presence at both endpoints, and market-specific operational characteristics (such as total distance, metal carriers, and connection requirements) that maintain shift supply conditions while remaining plausibly independent of route-specific demand fluctuations. Similarly, demand-shifters include the total number of destinations accessible from the origin airport (excluding merging parties’ routes) and demographic variables (population-based market size, births, deaths, net migration, total income, and per capita income) that influence passenger demand without directly affecting operational costs or strategic pricing considerations.
4.2.2 Strategic Conduct Parameters
Lemma 1 establishes identification of all demand parameters and the nuisance parameters $\tilde{\gamma}_{0}$ and $\tilde{\gamma}_{1}$. However, because these parameters conflate three distinct model primitives: the economies of scale ($\theta_{1}$), and the pre-merger conduct ($\lambda_{\text{pre}}=\lambda(I_{mt}=0)$) and post-merger conduct ($\lambda_{\text{post}}=\lambda(I_{mt}=1)$) parameters, we cannot separately identify these three primitives. However, we can identify the merger-induced change in strategic conduct, i.e., $\Delta\lambda=\lambda_{\text{post}}-\lambda_{\text{pre}}$. More formally show this result in Lemma 2-1.
To identify these three parameters, we require additional information. While we do not use this identification strategy in our empirical analysis, for completeness of the argument, we note that one potential source of such information could come from heterogeneity in demand elasticity across time or market segments—for instance, due to varying consumer compositions (business or leisure travelers) during peak versus off-peak travel periods or across different routes (such as holiday or business routes). In Lemma 2-2 we show that if we have two demand slopes, it would be sufficient to identify $\theta_{1}$, $\lambda_{\text{pre}}$, and $\lambda_{\text{post}}$.
Lemma 2.
Suppose all the assumptions in Lemma 1 hold.
  1. 1.
    The supply and strategic conduct parameters ($\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}}$) cannot be individually identified. However, the change in strategic conduct, defined as $\Delta\lambda:=\lambda_{\text{post}}-\lambda_{\text{pre}}$, can be identified as $\Delta\lambda=-\alpha_{1}\times(\tilde{\gamma}_{1}-\tilde{\gamma}_{0})$.
  2. 2.
    Suppose there are two slope of the demand function $(\alpha_{1}^{1},\alpha_{1}^{2})$. Then, the supply and conduct parameters ($\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}}$) are identified.
The proof is in Appendix B, but here we provide an intuitive explanation of the results. The non-identification in Lemma 2-(1) arises from the substitutability between marginal cost and conduct parameters in determining $\tilde{\gamma}_{mt}$. For illustration, suppose we increase the strategic conduct, i.e., make firms less competitive, by $\kappa>0$ and the scale parameter $\theta_{1}$ by $\kappa/\alpha_{1}$. On the one hand, the economies of scale put downward pressure on prices. On the other hand, because firms are now less competitive, there is an upward pressure on prices. In the proof, we show that these opposing pressures on prices cancel out, i.e., the two sets of parameters $(\lambda_{\text{pre}},\lambda_{\text{post}},\theta_{1})$ and $(\kappa\lambda_{\text{pre}},\kappa\lambda_{\text{post}},\theta_{1}\kappa/\alpha_{1})$ are observationally equivalent. However, the change in the strategic conduct ($\Delta\lambda$) is identified as $\Delta\lambda=-\alpha_{1}\times(\tilde{\gamma}_{1}-\tilde{\gamma}_{0})$.
Lemma 2-(2) shows that if we have two demand functions with different slopes $(\alpha_{1}^{1}\neq\alpha_{1}^{2})$, then we can separate strategic conduct from the economies of scale parameter by breaking the above-mentioned substitutability between marginal costs and conduct. To illustrate this mechanism, suppose ${\mathcal{R}}_{mt}\in\{0,1\}$ that captures different demand regimes to (7), which gives
$Q_{mt}$ $=\alpha_{0}+\alpha_{1}^{1}P_{mt}+\alpha_{1}^{2}(P_{mt}\times{\mathcal{R}}_{mt})+\alpha_{2}Z_{mt}+\varphi_{mt},$
$P_{mt}$ $=\theta_{0}+\tilde{\gamma}_{mt}^{\text{new}}Q_{mt}+\theta_{2}W_{mt}+\theta_{3}I_{mt}+\vartheta_{mt},$
where $\tilde{\gamma}_{mt}^{\text{new}}=\theta_{1}-\frac{\lambda(I_{mt})}{\alpha_{1}^{1}+\alpha_{1}^{2}\times{\mathcal{R}}_{mt}}$. The crucial feature is that ${\mathcal{R}}_{mt}$ satisfies an exclusion restriction: it affects demand through the price elasticity but does not directly affect the supply. This exclusion restriction generates differential responses to identical conduct parameters across demand regimes. The conduct enters as $\lambda(I_{mt})/(\alpha_{1}^{1}+\alpha_{1}^{2}\times{\mathcal{R}}_{mt})$, the same underlying strategic behavior manifests differently when ${\mathcal{R}}=0$ versus ${\mathcal{R}}=1$, providing the variation needed to separate conduct from costs. The rank condition also confirms this argument. With two demand slopes, there are four coefficients of $Q$, i.e., $\tilde{\gamma}_{mt}^{\text{new}}$ that, after some simplification, are given by the following system of equations
$\begin{bmatrix}\tilde{\gamma}_{0}^{1}\\ \tilde{\gamma}_{1}^{1}\\ \tilde{\gamma}_{0}^{2}\\ \tilde{\gamma}_{1}^{2}\end{bmatrix}=\begin{bmatrix}1&-\frac{1}{\alpha_{1}^{1}}&0\\ 1&0&-\frac{1}{\alpha_{1}^{1}}\\ 1&-\frac{1}{\alpha_{1}^{2}}&0\\ 1&0&-\frac{1}{\alpha_{1}^{2}}\end{bmatrix}\begin{bmatrix}\theta_{1}&\lambda_{\text{pre}}&\lambda_{\text{post}}\end{bmatrix}^{\top}$.
As $\alpha_{1}^{1}\neq\alpha_{1}^{2}$, the coefficient matrix has full column rank, identifying $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{pre}})$.
4.3 First Bridge: Structural to Quasi-Experimental
We now demonstrate how we can bridge between our structural framework and the DiD approach. The key insight is that our structural equation model of demand and supply generates a reduced-form equation with the same structure as the DiD specification. In particular, by substituting the demand into the supply equation in Equation (7), we get the following “reduced-form” price equation:
$P_{mt}=\frac{\theta_{0}+\tilde{\gamma}_{mt}\alpha_{0}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}+\frac{\theta_{3}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}I_{mt}+\frac{\tilde{\gamma}_{mt}\alpha_{2}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}Z_{mt}+\frac{\theta_{2}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}W_{mt}+\frac{\tilde{\gamma}_{mt}\varphi_{mt}+\vartheta_{mt}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}.$ (9)
This equation maps directly to the DiD estimation Equation (1). The intercept corresponds to the combined effect of constants plus market and time fixed effects, the exogenous variables $X_{mt}$ represent the demand and cost shifters $(Z_{mt},W_{mt})$, and $I_{mt}$ is the treatment indicator. [^9] The dependent variables in Equation (1) are log prices, while they are (level) prices in Equation (9).
4.3.1 Aggregate Price Effect of Merger
An application of this equivalence is that researchers seeking an aggregate measure of merger effects on prices can still use our structural framework. In particular, the structural approach recovers the same interpretable coefficient while being robust to the fundamental identification challenges that may render the standard DiD estimate unreliable.
In particular, comparing Equations (1) and (9), the coefficient on the merger indicator $I_{mt}$ can be thought of as the aggregate price effect of the merger, which is:
$\beta^{\text{DiD}}=\frac{\theta_{3}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}=\frac{\theta_{3}}{1-\frac{{\mathcal{E}}_{d,mt}}{{\mathcal{E}}_{s,mt}}}=\frac{\text{cost efficiency}}{\text{cost pass-through}},$ (10)
where the second last equality uses the fact that the demand elasticity is ${\mathcal{E}}_{d,mt}=\alpha_{1}\frac{P_{mt}}{Q_{mt}}$, the supply elasticity is ${\mathcal{E}}_{s,mt}=\frac{1}{\tilde{\gamma}_{mt}}\frac{P_{mt}}{Q_{mt}}$, and their ratio is $\tilde{\gamma}_{mt}\alpha_{1}$, and the last equality uses the definition of cost pass-through rate (Weyl and Fabinger, 2013). [^10] For more on merger effects and pass-through rates, see Jeffe and Weyl (2013) and Miller et al. (2016). Equation (10) clarifies that the aggregate effect of the merger equals the cost efficiency adjusted for the cost pass-through rate. It captures if and how merger-induced cost changes are passed to prices, given the market’s demand and supply elasticities.
Any such aggregate measure conflates the direct cost effect ($\theta_{3}$) with conduct changes embedded in $\tilde{\gamma}_{mt}$. A small coefficient could indicate either modest efficiency gains or low pass-through due to increased market power. Our structural approach resolves this ambiguity by identifying $\theta_{3}$, the actual cost efficiency, separate from the pass-through rate determined by demand and supply estimates, which includes the firms’ strategic conduct. As we demonstrate next, when mergers affect strategic conduct and alter supply elasticity, our structural framework can detect and correct the resulting biases that affect DiD design.
4.3.2 Robustness
To understand why structural estimates are robust while DiD estimates may not be, we examine the error structure revealed by the reduced-form Equation (9). The error in DiD is not a simple exogenous shock but rather a composite
$\varepsilon_{mt}=\frac{\tilde{\gamma}_{mt}\varphi_{mt}+\vartheta_{mt}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}=\frac{\tilde{\gamma}_{mt}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}\varphi_{mt}+\frac{1}{(1-\tilde{\gamma}_{mt}\alpha_{1})}\vartheta_{mt}:=A_{mt}\varphi_{mt}+B_{mt}\vartheta_{mt},$
of demand shocks $\varphi_{mt}$ and supply shocks $\vartheta_{mt}$, where the weights $A_{mt}$ and $B_{mt}$ depend on model primitives including the strategic conduct parameter embedded in $\tilde{\gamma}_{mt}$; see Equation (10). Because mergers can alter strategic conduct, they potentially change these very weights that determine how structural shocks manifest in observed prices, generating endogeneity that the DiD design cannot remedy.
Standard parallel trends assumptions must hold for this combined error, not for the underlying economic shocks. This requirement is more restrictive than typically acknowledged and breaks down precisely when mergers have their most significant effects: altering firms’ strategic conduct. We illustrate this through two scenarios.
Case 1: Constant Strategic Conduct. When strategic conduct remains unchanged following a merger, the weights $A_{mt}$ and $B_{mt}$ are invariant to treatment status. The parallel trends requirement becomes:
$\mathbb{E}[\varepsilon_{m,\text{post}}-\varepsilon_{m,\text{pre}}|I_{mt}=1]=\mathbb{E}[\varepsilon_{m,\text{post}}-\varepsilon_{m,\text{pre}}|I_{mt}=0].$ (11)
Substituting $\varepsilon_{mt}=A_{mt}\varphi_{mt}+B_{mt}\vartheta_{mt}$ requires parallel trends for demand and supply shocks:
$E[\varphi_{m,\text{post}}-\varphi_{m,\text{pre}}|\text{Treatment}=1]=E[\varphi_{m,\text{post}}-\varphi_{m,\text{pre}}|\text{Treatment}=0],$ (12)
$E[\vartheta_{m,\text{post}}-\vartheta_{m,\text{pre}}|\text{Treatment}=1]=E[\vartheta_{m,\text{post}}-\vartheta_{m,\text{pre}}|\text{Treatment}=0].$ (13)
Airlines systematically select markets based on both demand prospects and cost considerations, violating at least one of these conditions. The DiD approach requires both these conditions to be satisfied, whereas the structural approach requires only the supply-side parallel trends condition (13) to be satisfied. Selection on demand unobservables does not threaten identification because mergers affect costs directly, not demand, and we instrument for endogenous quantities using our exclusion restrictions.
Case 2: Varying Strategic Conduct. When mergers alter strategic conduct, so $\lambda_{\text{post}}>\lambda_{\text{pre}}$, which in turn means $\tilde{\gamma}_{m,\text{post}}>\tilde{\gamma}_{m,\text{pre}}$ for treated markets while $\tilde{\gamma}_{m,\text{post}}=\tilde{\gamma}_{m,\text{pre}}$ for control markets, the equilibrium weights $A_{mt}$ and $B_{mt}$ become endogenous to the merger decision. The change in the error term for treated markets becomes: $\varepsilon_{m,\text{post}}-\varepsilon_{m,\text{pre}}=A_{m,\text{post}}\varphi_{m,\text{post}}-A_{m,\text{pre}}\varphi_{m,\text{pre}}+B_{m,\text{post}}\vartheta_{m,\text{post}}-B_{m,\text{pre}}\vartheta_{m,\text{pre}}.$ Since weights change for treated markets but remain constant for control markets, the parallel trends condition in Equation (11) is mechanically violated even when both underlying structural shocks satisfy perfect parallel trends. This feature represents a breakdown of DiD identification due to simultaneity between demand and supply.
When mergers increase market power, the weight $A_{mt}$ on demand shocks increases. Any positive demand shock in the post-merger period generates an amplified price response through enhanced market power. The DiD specification attributes this entire price increase to the merger dummy, confounding the direct merger effect with the interaction between demand shocks and changed conduct. This creates systematic bias that increases with both demand volatility and conduct changes.
DiD approaches to merger evaluation are therefore vulnerable to endogeneity problems extending beyond standard parallel trends concerns. The composite error structure creates mechanical correlations between treatment and unobservables whenever mergers affect the economic mechanisms that determine equilibrium outcomes. Our structural methodology avoids these limitations by explicitly modeling the conduct parameter and using different identifying variation. By separating efficiency effects from conduct changes, the structural approach eliminates the mechanical correlation between the error term and treatment status.
5 Estimation of the Structural Model
Having established the framework and identification strategy, we now turn to estimating the structural model. This section presents estimates of the key parameters that generate the merger effects. We focus our presentation on the parameters most critical for merger evaluation—merger-induced efficiency, changes in strategic conduct, and demand price elasticity—estimated using the first differences method. We employ several instruments for each Equation in the simultaneous system (7) to strengthen identification relative to the stylized single-instrument case presented in the previous section, which serves for theoretical clarity.
5.1 Instrumental Variables
Our identification strategy leverages two distinct sources of exogenous variation: demand shifters that affect consumer travel decisions without entering airline cost functions, and cost shifters that influence supply conditions without directly affecting consumer utility.
5.1.1 Demand Shifters
We employ two complementary sets of demand instruments that exploit different sources of variation in the extensive margin decision to fly.
Our primary demand instrument exploits variation in the total number of markets served from origin airports, excluding routes operated by merging airlines. This measure captures the option value of network access: broader destination availability increases travel flexibility, enables multi-city trips, and raises the probability that air travel meets specific transportation needs. Crucially, this market thickness effect makes air travel more attractive relative to driving, without affecting any individual airline’s operational costs. While carrier-specific networks affect marginal costs (Ciliberto et al., 2021), the aggregate count of destinations served by all carriers affects only demand through increased consumer awareness and consideration of air travel. We also include population-based market size (geometric mean).
5.1.2 Cost Shifters
We utilize two categories of supply-side instruments that impact competitive conditions and operational costs without directly affecting consumer preferences.
We measure the number of airlines operating at origin and destination airports, capturing the pool of potential entrants. Following Goolsbee and Syverson (2008), carrier presence at both endpoints reduces entry costs and intensifies competitive pressure on pricing and entry decisions. We separately instrument using legacy carrier presence, as these airlines face different costs and competitive dynamics than low-cost carriers. We also include a categorical variable to control for whether Southwest is a potential entrant.
We also count the number of distinct regional carriers used by a legacy carrier to serve a specific market. The use of regional airlines should be correlated with costs (through fuel consumption, crew requirements, aircraft utilization, and network management). However, it should not directly impact consumer utility, given the observed product characteristics.
These instruments satisfy the exclusion restrictions necessary for identification and provide substantial variation for estimating the simultaneous equations system.
5.2 Estimated Merger Effects
Consistent with our DiD analysis, we allow efficiency effects to materialize over the two years following a merger and use our preferred first-difference method for estimation. The estimation equation for prices in (7) becomes:
$P_{mt}=\theta_{0}+\left(\theta_{1}-\frac{\lambda_{mt}}{\alpha_{1}}\right)\times Q_{mt}+\theta_{2}\times W_{mt}+\sum_{k=1}^{8}\theta_{3,k}\times I_{m,t,k}+\text{Market FE}+\vartheta_{mt}$ (14)
where $I_{m,t,k}=1$ if market $m$ is in quarter $k$ post-merger at time $t$, and 0 otherwise. This estimation recognizes that mergers may facilitate coordination among remaining competitors. As before, we continue to assume that the efficiency may take time to realize, but we assume that the firms’ strategic conduct may change immediately after the merger. We allow the conduct parameter to vary across periods as:
[…]
Correspondingly, let $D^{\text{ctrl}}_{mt}$ be a binary variable equal to one if $m$ is a control market at time $t$, and zero otherwise. Similarly, let $D^{\text{treat-pre}}_{mt}$ and $D^{\text{treat-post}}_{mt}$ be equal to one if $m$ is treated and pre-merger (respectively, post-merger) at time $t$, and zero otherwise. Finally, let $\text{Transition}_{mt}$ be equal to one only when a market $m$ transitions from pre- to post-merger, and $\text{Transition}_{m,t,k}$ be equal to one when transitioning specifically into quarter $k$ post-merger. So, we keep track of three cases: $\text{Cases}=\{\text{ctrl, treat-pre, treat-post}\}$. Then, we can write the estimating price equation in first difference as
$\Delta P_{mt}=$ $\left[\sum_{j\in\text{Cases}}\left(\theta_{1}-\frac{\lambda_{j}}{\alpha_{1}}\right)\times D^{j}_{mt}\right]\times\Delta Q_{mt}-\frac{\Delta\lambda}{\alpha_{1}}\times Q_{m,t-1}\times\text{Transition}_{mt}$ (15)
$+\theta_{2}\times\Delta W_{mt}+\sum_{k=1}^{8}\theta_{3,k}\times\text{Transition}_{m,t,k}+\Delta\vartheta_{mt}$
The key insight is that the slope coefficient on $\Delta Q_{mt}$ differs across market types and time periods, reflecting different conduct parameters. Additionally, there is a level effect involving $Q_{m,t-1}$ when markets transition from a pre-merger to a post-merger state due to changes in conduct parameters. Each coefficient $\theta_{3,k}$ represents the price change when transitioning into quarter $k$ after the merger. The average efficiency effect weights each quarter’s contribution by the fraction of post-merger observations that experience it: $\sum_{k=1}^{8}\theta_{3,k}\times\frac{9-k}{8}$. Table 5 presents the estimation results.
Demand Parameters. Panel A shows the estimated demand slopes, which are consistently negative and statistically significant across all three mergers: -1.640 for DL-NW, -1.761 for UA-CO, and -3.112 for AA-US. These coefficients translate into median demand elasticities ranging from -0.884 to -1.944 for control markets, and -0.804 to 1.609 for treated markets, indicating that market-level airline demand is relatively inelastic.
Table 5: Estimation Results
DL-NW UA-CO AA-US
Panel A
Demand Slope -1.640*** -1.761*** -3.112***
(0.259) (0.183) (0.322)
Demand Elasticity (Median)
 - Control Markets -0.884*** -0.981*** -1.944***
(0.139) (0.102) (0.201)
 - Treated Markets -1.609*** -0.804*** -1.492***
(0.254) (0.084) (0.154)
Panel B
Change in Conduct -0.141 0.031 0.107**
(0.101) (0.022) (0.052)
Panel C
Average Efficiency -0.119 0.030 0.140
(0.169) (0.113) (0.107)
Panel D
Aggregate Price Effect of Merger -0.086 0.026 0.106
(0.108) (0.099) (0.084)
Observations 13,304 12,299 12,458
  • Note: Table presents two-stage least squares estimates allowing for strategic conduct changes across pre- and post-merger periods for three airline mergers. Panel A shows demand slope and implied elasticities for control and treated markets. Panel B presents changes in strategic conduct parameter averaged over eight post-merger quarters, where positive values indicate increased coordination among remaining competitors. Panel C reports merger-induced average cost efficiency. Average efficiency is the weighted average of quarterly efficiency effects, where each quarter is weighted by the proportion of post-merger observations experiencing that effect. Panel D presents aggregate price effects of mergers. Price effects are calculated using Equation (10), where the numerator is the average efficiency from Panel C and the denominator incorporates the estimated demand and supply parameters. Standard errors in parentheses. Weak Instruments: demand equations achieve Kleibergen-Paap F-statistics of 24.59 (DL-NW), 47.98 (UA-CO) and 70.20 (AA-US), all exceeding the Stock-Yogo critical value of 24.58. The supply equations have Kleibergen-Paap F-statistics of 128.14 (DL-NW), 61.19 (UA-CO), and 123.55 (AA-US), all substantially exceeding the Stock-Yogo critical value of 22.30. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
Change in Strategic Conduct. Panel B reveals evidence of post-merger coordination effects that vary substantially across mergers. For DL-NW, we observe a change in conduct parameter of $\Delta\lambda=-0.141$, indicating a decrease in coordination following the merger, although this estimate is statistically insignificant. This result likely reflects the limited number of control markets available for this merger, as the DL-NW transaction involved two large network carriers with extensive route overlap, leaving fewer comparable control markets.
UA-CO shows a modest increase in the conduct parameter ($\Delta\lambda=0.031$), though this estimate is also imprecise. In contrast, AA-US demonstrates a more substantial increase ($\Delta\lambda=0.107$, p-value 0.05)—a meaningful shift toward less competitive behavior among the remaining airlines, providing evidence that mergers can have coordinated effects.
Average Efficiency. Panel C shows the effect of mergers on cost efficiency. DL-NW exhibits efficiency gain of $\mathdollar 11.9$ per hundred dollars of fare. UA-CO shows modest efficiency loss of $\mathdollar 3$ per hundred dollars, while AA-US demonstrates efficiency gains of $\mathdollar 14$ per hundred dollars. However, these efficiency parameters are imprecisely estimated, so we should exercise caution in drawing conclusions about efficiency and mergers.
Aggregate Price Effects. Panel D presents the aggregate merger effects on prices, incorporating both efficiency and conduct channels. As with the previous estimates, none of the aggregate price effects is estimated precisely. DL-NW shows a price decrease of $\mathdollar 8.6$ per hundred dollars of fare, reflecting the combination of efficiency gain and reduced conduct. UA-CO exhibits a modest price increase of $\mathdollar 2.6$ per hundred dollars, while AA-US shows a larger price increase of $\mathdollar 10.6$ per hundred dollars. These effects reflect the combined impact of efficiency changes and altered competitive conduct following the merger.
The progressive deterioration in competitive conduct – from negative to positive to statistically significant positive – suggests that successive mergers may have facilitated increasingly coordinated behavior among the remaining carriers. Indeed, this escalating pattern of coordination effects was precisely the concern that led the US DOJ to initially block the AA-US merger in 2013, fearing that the merger would generate coordinated effects. While we cannot make definitive causal claims due to statistical limitations, the systematic temporal progression provides suggestive evidence that merger waves in concentrated industries may have cumulative effects on market competitiveness.
5.3 Second Bridge: Quasi-Experimental to Structural
Our merger analysis faces a challenge: the small number of treated markets relative to controls makes it difficult to obtain meaningful and precise parameter estimates. The three mergers, DL-NW, UA-CO, and AA-US, 14 out of 836, 6 out of 774, and 12 out of 779 markets, respectively, each representing a small fraction of the total route network. This limited treatment variation may affect the estimates of efficiency gains and changes in conduct, as well as their precision.
Selecting suitable control markets for merger analysis is a challenging task. The ideal control group should closely mirror the counterfactual evolution of treated markets in the absence of the merger, thereby satisfying the parallel trends assumption that is crucial for DiD estimation. As Ashenfelter et al. (2009) observe, researchers face an inherent tradeoff: control markets must be geographically distant enough to avoid merger spillovers yet similar enough to face comparable demand and cost conditions. This tradeoff presents what Ashenfelter et al. (2009) term the “similarity-independence tension” in control market selection.
Our use of first-difference methods compounds the estimation challenge. Griliches and Mairesse (1999) demonstrate that first-difference estimators, while addressing fixed effects, can substantially reduce identifying variation. Our results confirm this pattern: first-difference specifications consistently yield larger standard errors and less significant coefficients compared to fixed effects approaches (Table C.2). However, fixed effects estimates may be biased due to unobserved market characteristics that are correlated with the merger treatment.
We propose a solution that, in keeping with the overarching goal of the paper, bridges the quasi-experimental literature with structural econometrics. Our approach, which we term synthetic GMM, combines attractive features of synthetic control methods (Abadie and Gardeazabal, 2003) and synthetic DiD (Arkhangelsky et al., 2021) with structural estimation. Like synthetic DiD, our method reweights control markets using pre-treatment price similarity to construct better counterfactuals, thereby strengthening identification. The key innovation is straightforward and tractable: we utilize the market- and time-specific synthetic DiD weights to reweight each market’s contribution to the moment conditions.
Table 6: Algorithm for Synthetic GMM
1:Estimation:
2:Data: $\{P_{mt},(\boldsymbol{W}_{mt},\boldsymbol{Z}_{mt},Q_{mt})\}$ with treated units $\mathcal{T}$ and control units $\mathcal{C}$
3:Construct residualized pre-merger outcome (prices): $\bar{P}_{tr,pre}$ (treated averages), and $P_{co,pre}$ (control matrix)
4:Solve for market weights: $\hat{\omega}=\arg\min_{\omega}\|\bar{P}_{tr,pre}-\omega^{\prime}P_{co,pre}\|_{2}^{2}+\zeta\|\omega\|_{2}^{2}$
5:Solve for time weights: $\hat{\tau}=\arg\min_{\tau}\|\bar{P}_{tr,pre}^{\prime}-\tau^{\top}P_{tr,pre}^{\top}\|_{2}^{2}+\zeta\|\tau\|_{2}^{2}$
6:Define observation weights: $\omega_{mt}=\hat{\omega}_{m}$ if $m\in\mathcal{C}$, else 11; $r_{mt}=\hat{\tau}_{t}$ if pre-treatment, else 11
7:Synthetic moments: $\tilde{g}(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda})=\frac{1}{MT}\sum_{i,t}\omega_{mt}\cdot r_{mt}\cdot g(P_{mt},\boldsymbol{W}_{mt},\boldsymbol{Z}_{mt},Q_{mt};(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))$
8:Synthetic GMM: $(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}}):=\arg\min_{(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda})}\tilde{g}((\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))^{\top}\hat{W}\tilde{g}((\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))$
9:Bootstrap Inference:
10:for $b=1,\ldots,B$ do
11:  Resample units with replacement (preserving time series and treatment status)
12:  Estimate weights $\hat{\omega}^{(b)},\hat{\tau}^{(b)}$ for bootstrap sample
13:  Estimate synthetic GMM $(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}})^{(b)}$ using bootstrap weights
14:end for
15:Compute bootstrap standard errors from $\{(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}})^{(b)}\}_{b=1}^{B}$
More formally, our synthetic GMM approach proceeds in two main steps, as shown in the Table 6. First, following Arkhangelsky et al. (2021), we construct market-specific weights $\hat{\omega}$ that align pre-merger price trends between control and treated markets, and time-specific weights $\hat{\tau}$ that balance pre-merger periods. Like synthetic control methods, we find weights such that the weighted average of control market residualized prices closely matches treated market residualized prices: $\sum_{m\in\text{control}}\hat{\omega}_{m}P_{mt}\approx\frac{\sum_{m\in\text{treated}}P_{mt}}{N_{\text{treated}}}$ across all pre-merger periods $t$. The time weights similarly balance temporal variation between the two groups.
Second, we integrate these weights directly into our GMM estimation by reweighting each moment condition, which we refer to as synthetic GMM. Instead of the standard sample moments $\frac{1}{MT}\sum_{m,t}g(\cdot;\cdot)$, we compute synthetic-weighted moments $\frac{1}{MT}\sum_{m,t}\omega_{mt}r_{mt}g(\cdot;\cdot)$, where $\omega_{mt}$ applies market weights to control units and $r_{mt}$ applies time weights to pre-treatment periods. This reweighting ensures that parameter identification emphasizes the most informative control observations while preserving the structural interpretation of synthetic GMM estimates. For inference, we implement block bootstrap resampling that jointly estimates weights and structural parameters, accounting for uncertainty in both components.
Table 7: Estimated Merger Effects (Synthetic GMM)
DL-NW UA-CO AA-US
Panel A
Change in Conduct -0.001 0.049* 0.100**
(0.010) (0.027) (0.039)
Panel B
Average Efficiency 0.088 0.135* 0.134
(0.148) (0.074) (0.086)
Panel C
Aggregate Price Effect of Merger 0.085 0.129* 0.116
(0.140) (0.070) (0.075)
Observations 13,304 12,299 12,458
  • Note: Table presents weighted first differences two-stage least squares estimates for three airline mergers. Panel A presents changes in strategic conduct parameter averaged over eight post-merger quarters, where positive values indicate increased coordination among remaining competitors. Panel B shows average merger-induced cost efficiency over the same period. Panel C reports aggregate price effects incorporating both efficiency and conduct channels. Estimation uses market-size weights. Standard errors in parentheses. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
Table 7 displays the estimates. The estimates suggest heterogeneity in merger effects across the three airline consolidations, with notably improved precision compared to the estimates in Table 5. The weighting approach yields smaller standard errors across all parameters, while maintaining the same qualitative conclusions about merger impacts. The estimates for the UA-CO merger suggest increased strategic conduct among remaining competitors by 0.049 (p-value 0.1), suggesting enhanced coordination in the post-merger market. Simultaneously, the merger appears to have reduced efficiency, i.e., increased cost, by \$13.5, which represents approximately 5.9% of the average pre-treatment price of \$228.25.
The AA-US merger shows a different but equally troubling pattern. The conduct parameter increased by 0.100 (p-value 0.05), representing a substantial shift toward more coordinated behavior among competitors. For interpretability, suppose that pre-merger conduct was around 0.2 (reflecting moderate competition), an increase of 0.100 represents a 50% rise in the conduct parameter, moving the market significantly closer to joint profit maximization. The UA-CO increase of 0.049 would represent a 25% increase from the same baseline, still indicating meaningful coordination effects.
These findings underscore the importance of rigorous merger analysis using a structural approach. The synthetic GMM approach can improve our ability to detect conduct changes that might be missed with standard GMM. The results suggest that airline mergers were not pro-competitive as they generated limited efficiency gains while facilitating coordination.
We conclude with the following observation. A natural concern with our synthetic GMM approach is whether the reweighting procedure, combined with first-difference estimation, still eliminates too much identifying variation—potentially explaining why several parameters remain statistically insignificant. To address this concern, we conducted a robustness check by estimating the identical synthetic GMM model using fixed effects.
The fixed effects results, detailed in the appendix, initially validate merger efficiency claims: all three consolidations show statistically significant efficiency gains and pro-competitive effects on pricing. However, when we use synthetic GMM—creating an “apples-to-apples” comparison—these apparently favorable merger effects largely disappear, yielding estimates remarkably similar to those from our first-difference results. The similarity between synthetic-weighted first differences and fixed effects is reassuring that our approach captures genuine effects.
In summary, the DL-NW merger has a minimal impact across all dimensions, consistent with a consolidation that neither significantly enhances efficiency nor facilitates coordination. The UA-CO and AA-US mergers present concerning patterns. The UA-CO merger increased competitor coordination and reduced efficiency, resulting in approximately 5% higher prices. The AA-US merger, which occurred last in the sequence, exhibits the largest coordination effects, but because the merger-induced efficiency is imprecisely estimated, price changes are also statistically insignificant.
6 Conclusion
This paper establishes a connection between quasi-experimental and structural approaches to merger evaluation. We build a two-way methodological bridge. The first bridge shows that the widely used DiD estimation equation emerges as the reduced form of a structural demand and supply equation, which in turn clarifies the reasons why DiD identification may fail in the merger context. The second bridge demonstrates that we can leverage recent advancements in quasi-experimental approaches, particularly synthetic control methods, to enhance the estimation of structural models (synthetic GMM). This bridge reveals that these seemingly distinct methodologies represent different levels of economic detail applied to the same underlying relationships, and offers a unified framework that combines the interpretability of DiD with the robustness of structural identification.
We demonstrate that the DiD coefficient represents a composite parameter conflating cost efficiencies, economies of scale, and strategic conduct changes—effects that can only be separated through structural decomposition. Consequently, mergers present fundamental identification challenges for DiD approaches: when mergers alter strategic behavior, the error structure in DiD specifications becomes endogenous to treatment status, resulting in systematic bias. Our structural framework addresses these limitations while enabling researchers to construct DiD-equivalent aggregate measures from structural parameters, obtaining both methodological robustness and economic interpretability.
Our empirical analysis of three major airline mergers reveals that apparent price stability masks offsetting efficiency and conduct effects. Under the assumptions of perfect competition, we find minimal efficiency gains across all mergers. Allowing for strategic conduct changes meaningful coordination effects among remaining competitors for UA-CO and AA-US. These findings suggest that modest efficiency gains are largely offset by increased strategic coordination, consistent with the industry exhibiting features that facilitate coordination (Ciliberto and Williams, 2014; Aryal et al., 2022). The structural decomposition provides crucial insights that pure reduced-form analysis cannot capture, demonstrating why merger evaluation requires attention to underlying economic mechanisms.
Our methodological innovations have broader implications beyond our specific application. The sensitivity of DiD estimates to specification choices raises questions about existing merger retrospectives that rely solely on quasi-experimental methods. Our structural approach, while requiring only standard instrumental variables combined with synthetic control reweighting, offers a path toward a more robust merger evaluation that captures the underlying economic mechanisms.
Although focused on airlines, our bridging approach offers broad applicability for analyzing mergers in concentrated industries where strategic conduct plays a central role. The methodological framework provides a foundation for robust merger retrospectives that can inform both academic research and antitrust policy (Asker and Nocke, 2021).
We maintain homogeneity assumptions and employ market-level analysis, which enables direct comparison with quasi-experimental approaches. Looking ahead, promising research directions include adapting our bridging methodology to incorporate product differentiation with a nonlinear demand system (Nevo, 1998; Berry et al., 1995; Berry and Haile, 2014).
Appendix
A Data Construction
A.1 Origin and Destination Survey (DB1B)
Our primary data source is the Domestic Origin and Destination Survey (DB1B), a 10% sample of airline tickets from all reporting carriers. This dataset is available from the U.S. DOT’s National Transportation Library. We clean the data by removing the following:
  • Tickets with more than six coupons or more than three coupons in either direction.
  • Tickets sold by foreign carriers flying between two U.S. airports.
  • Tickets that are part of international travel.
  • Tickets involving non-contiguous domestic travel (Hawaii, Alaska, and Territories).
  • Tickets with questionable fare credibility or bulk fare indicators.
  • Tickets that are neither one-way nor round-trip.
  • Interline tickets (travel on more than one airline in a directional trip).
We treat round-trip tickets as two separate directional trips, with the price for each direction being half of the round-trip price. A passenger flying on a round-trip ticket is counted as a single passenger in both directions of travel. We exclude tickets cheaper than \$25, those in the top and bottom 1% of the year-quarter fare distribution, and airlines serving fewer than 90 passengers per quarter.
A.2 T-100 Domestic Market (U.S. Carriers)
This dataset contains domestic nonstop segment data reported by U.S. air carriers, including carrier, origin, destination, and available capacity. We use this dataset to obtain information on carriers’ capacity on their nonstop routes.
A.3 Additional Data Tables
Table A.1: Summary Statistics
Variable Mean Median SD
DL-NW
Quantity (10,000s) 3.4041 2.5625 3.1592
Price (\$100) 1.8717 1.8349 0.5074
Market size (1,000,000s) 3.6211 3.1279 1.9928
Per-Capita Income at Origin (\$10,000) 42.0451 41.0382 6.5522
Per-Capita Income at Destination (\$10,000) 42.0489 41.0564 6.5502
Net Migration at Origin (100) 0.0979 0.0515 0.3162
Net Migration at Destination (100) 0.0975 0.0515 0.3157
Population at Origin (1,000,000) 4.8210 3.1827 4.6315
Population at Destination (1,000,000) 4.8094 3.1032 4.6361
Births in Market (100,000) 22,242.79 19,689.00 16,731.96
Deaths in Market (100,000) 10,694.55 8,820.00 8,073.81
Ratio of Nonstop passengers 0.8409 0.8937 0.1779
UA-CO
Quantity (10,000s) 3.4753 2.5430 3.4403
Price (\$100) 1.9410 1.8765 0.5183
Market size (1,000,000) 3.7576 3.2795 2.0466
Per-Capita Income at Origin (\$10,000) 43.7443 42.9350 7.1511
Per-Capita Income at Destination (\$10,000) 43.7459 42.9350 7.1509
Net Migration at Origin (100) 0.1432 0.0596 0.2420
Net Migration at Destination (100) 0.1432 0.0596 0.2416
Population at Origin (1,000,000) 5.0334 3.3734 4.7191
Population at Destination(1,000,000) 5.0202 3.3400 4.7316
Births in Market (100,000) 23,283.55 20,683.20 15,904.49
Deaths in Market (100,000) 11,701.28 10,380.60 8,047.72
Ratio of Nonstop passengers 0.8484 0.8878 0.1518
AA-US
Quantity (10,000s) 3.3443 2.5250 3.1662
Price (100) 2.0874 2.0239 0.5027
Market size (1,000,000) 3.8320 3.4570 1.8798
Per-Capita Income at Origin (\$10,000) 48.6923 47.5981 7.9698
Per-Capita Income at Destination (\$10,000) 48.6793 47.5981 7.9674
Net Migration at Origin (100) 0.1941 0.1378 0.3408
Net Migration at Destination (100) 0.1935 0.1378 0.3400
Population at Origin (1,000,000) 5.2040 3.4762 4.8102
Population at Destination (1,000,000) 5.1877 3.4762 4.8144
Births in Market (100,000) 26,604.41 24,094.80 13,557.27
Deaths in Market (100,000) 14,331.84 12,780.80 7,384.71
Ratio of Nonstop passengers 0.8248 0.8688 0.1704
  • Note: Summary statistics table of the variables used to estimate the structural model. For each merger, the sample includes all pairwise combination of the top 50 populated origin and destination markets, restricted to three years leading up to the said merger.
Table A.1 presents summary statistics for the key variables employed in our empirical analysis. These variables capture various product characteristics, including price, service type, and carrier attributes that influence consumer choices in the airline industry. For each merger, the sample for these tables includes all pairwise combinations of the top 50 populated origin and destination markets, restricted to three years before the merger. This sampling approach ensures sufficient market coverage while maintaining a manageable dataset focused on the pre-merger competitive landscape.
B Proofs of the Identification Results
Lemma 1.
Consider the structural system governing equilibrium outcomes in market $m$ during period $t$:
$Q_{mt}$ $=\alpha_{0}+\alpha_{1}P_{mt}+\alpha_{2}^{\prime}\mathbf{Z}_{mt}+\varphi_{mt}\quad\text{(Demand)}$ (B.1)
$P_{mt}$ $=\theta_{0}+\tilde{\gamma}_{mt}Q_{mt}+\theta_{2}^{\prime}\mathbf{W}_{mt}+\theta_{3}I_{mt}+\vartheta_{mt}\quad\text{(Supply)}$ (B.2)
where $\mathbf{Z}_{mt}$ is a $d_{z}\times 1$ vector of demand shifters, $\mathbf{W}_{mt}$ is a $d_{w}\times 1$ vector of cost shifters, $I_{mt}$ is a scalar merger indicator, and $\varphi_{mt}$ and $\vartheta_{mt}$ represent unobserved demand and supply-side shocks, respectively. The parameters $\boldsymbol{\alpha}$ and $\boldsymbol{\theta}$ are $(2+d_{z})\times 1$ demand and supply parameter vectors, respectively.
The set of orthogonality conditions that ensure our instrumental variables provide exogenous variation is:
$\mathbb{E}[\mathbf{W}_{mt}\varphi_{mt}]$ $=\mathbf{0}_{d_{w}\times 1}\quad\text{(Cost shifters orthogonal to demand shocks)}$ (B.3)
$\mathbb{E}[\mathbf{Z}_{mt}\vartheta_{mt}]$ $=\mathbf{0}_{d_{z}\times 1}\quad\text{(Demand shifters orthogonal to supply shocks)}$ (B.4)
$\mathbb{E}[I_{mt}\vartheta_{mt}]$ $=0\quad\text{(Merger indicator orthogonal to supply shocks)}$ (B.5)
$\mathbb{E}[\mathbf{Z}_{mt}\varphi_{mt}]$ $=\mathbf{0}_{d_{z}\times 1}\quad\text{(Demand shifters orthogonal to demand shocks)}$ (B.6)
$\mathbb{E}[\mathbf{W}_{mt}\vartheta_{mt}]$ $=\mathbf{0}_{d_{w}\times 1}\quad\text{(Cost shifters orthogonal to supply shocks)}$ (B.7)
From the orthogonality conditions (B.3) and (B.6), we can construct moment conditions for demand identification. The orthogonality of cost shifters to demand shocks yields:
$\mathbb{E}[\mathbf{W}_{mt}(Q_{mt}-\alpha_{0}-\alpha_{1}P_{mt}-\alpha_{2}^{\prime}\mathbf{Z}_{mt})]=\mathbf{0}_{d_{w}\times 1}$ (B.8)
Additionally, the orthogonality of demand shifters to demand shocks provides:
$\mathbb{E}[\mathbf{Z}_{mt}(Q_{mt}-\alpha_{0}-\alpha_{1}P_{mt}-\alpha_{2}^{\prime}\mathbf{Z}_{mt})]=\mathbf{0}_{d_{z}\times 1}$ (B.9)
For supply identification, the orthogonality conditions (B.4), (B.5), and (B.7) generate the following moment conditions. The orthogonality of demand shifters to supply shocks gives:
$\mathbb{E}[\mathbf{Z}_{mt}(P_{mt}-\theta_{0}-\tilde{\gamma}_{mt}Q_{mt}-\theta_{2}^{\top}\mathbf{W}_{mt}-\theta_{3}I_{mt})]=\mathbf{0}_{d_{z}\times 1}$ (B.10)
The orthogonality of the merger indicator to supply shocks provides:
$\mathbb{E}[I_{mt}(P_{mt}-\theta_{0}-\tilde{\gamma}_{mt}Q_{mt}-\theta_{2}^{\top}\mathbf{W}_{mt}-\theta_{3}I_{mt})]=0$ (B.11)
Furthermore, the orthogonality of cost shifters to supply shocks yields:
$\mathbb{E}[\mathbf{W}_{mt}(P_{mt}-\theta_{0}-\tilde{\gamma}_{mt}Q_{mt}-\theta_{2}^{\top}\mathbf{W}_{mt}-\theta_{3}I_{mt})]=\mathbf{0}_{d_{w}\times 1}$ (B.12)
For notational simplicity, we define the following regressors in matrix form:
$\mathbf{X}_{D,mt}$ $=(1,P_{mt},\mathbf{Z}_{mt}^{\top})^{\top};\quad\!\!\mathbf{X}_{S,mt}=(1,Q_{mt},\mathbf{W}_{mt}^{\top},I_{mt})^{\top}$
$\mathbf{Z}_{D,mt}$ $=(1,\mathbf{W}_{mt}^{\top},\mathbf{Z}_{mt}^{\top})^{\top};\quad\!\!\mathbf{Z}_{S,mt}=(1,\mathbf{Z}_{mt}^{\top},I_{mt},\mathbf{W}_{mt}^{\top})^{\top}.$
The expanded set of moment conditions can then be expressed in matrix form. For demand parameters, combining equations (B.8) and (B.9) yields:
$\mathbb{E}[\mathbf{Z}_{D,mt}(Q_{mt}-\mathbf{X}_{D,mt}^{\top}\boldsymbol{\alpha})]=\mathbf{0}_{(1+d_{w}+d_{z})\times 1},$ (B.13)
which implies the system:
$\mathbb{E}[\mathbf{Z}_{D,mt}\mathbf{X}_{D,mt}^{\prime}]\boldsymbol{\alpha}=\mathbb{E}[\mathbf{Z}_{D,mt}Q_{mt}]$ (B.14)
For supply parameters, combining equations (B.10), (B.11), and (B.12) yields:
$\mathbb{E}[\mathbf{Z}_{S,mt}(P_{mt}-\mathbf{X}_{S,mt}^{\prime}\boldsymbol{\theta})]=\mathbf{0}_{(1+d_{z}+1+d_{w})\times 1}$ (B.15)
which corresponds to the system:
$\mathbb{E}[\mathbf{Z}_{S,mt}\mathbf{X}_{S,mt}^{\prime}]\boldsymbol{\theta}=\mathbb{E}[\mathbf{Z}_{S,mt}P_{mt}]$ (B.16)
For the demand parameters $\boldsymbol{\alpha}$ to be identified, the matrix $\mathbb{E}[\mathbf{Z}_{D,mt}\mathbf{X}_{D,mt}^{\prime}]$ must be of full column rank. This requires:
$\text{rank}(\mathbb{E}[\mathbf{Z}_{D,mt}\mathbf{X}_{D,mt}^{\prime}])=2+d_{z}$ (B.17)
Explicitly, this condition becomes:
$\text{rank}\begin{pmatrix}1&\mathbb{E}[P_{mt}]&\mathbb{E}[\mathbf{Z}_{mt}]^{\top}\\ \mathbb{E}[\mathbf{W}_{mt}]&\mathbb{E}[\mathbf{W}_{mt}P_{mt}]&\mathbb{E}[\mathbf{W}_{mt}\mathbf{Z}_{mt}^{\top}]\\ \mathbb{E}[\mathbf{Z}_{mt}]&\mathbb{E}[\mathbf{Z}_{mt}P_{mt}]&\mathbb{E}[\mathbf{Z}_{mt}\mathbf{Z}_{mt}^{\top}]\end{pmatrix}=2+d_{z}$ (B.18)
Similarly, for the supply parameters $\boldsymbol{\theta}$ to be identified, we require:
$\text{rank}(\mathbb{E}[\mathbf{Z}_{S,mt}\mathbf{X}_{S,mt}^{\top}])=3+d_{w}$ (B.19)
This condition translates to the explicit rank condition:
$\text{rank}\begin{pmatrix}1&\mathbb{E}[Q_{mt}]&\mathbb{E}[\mathbf{W}_{mt}]^{\top}&\mathbb{E}[I_{mt}]\\ \mathbb{E}[\mathbf{Z}_{mt}]&\mathbb{E}[\mathbf{Z}_{mt}Q_{mt}]&\mathbb{E}[\mathbf{Z}_{mt}\mathbf{W}_{mt}^{\top}]&\mathbb{E}[\mathbf{Z}_{mt}I_{mt}]\\ \mathbb{E}[I_{mt}]&\mathbb{E}[I_{mt}Q_{mt}]&\mathbb{E}[I_{mt}\mathbf{W}_{mt}]&\mathbb{E}[I_{mt}^{2}]\\ \mathbb{E}[\mathbf{W}_{mt}]&\mathbb{E}[\mathbf{W}_{mt}Q_{mt}]&\mathbb{E}[\mathbf{W}_{mt}\mathbf{W}_{mt}^{\top}]&\mathbb{E}[\mathbf{W}_{mt}I_{mt}]\end{pmatrix}=3+d_{w}$ (B.20)
The order conditions for identification are as follows. For the demand equation, we have $1+d_{w}+d_{z}$ instruments and $2+d_{z}$ parameters, requiring $1+d_{w}+d_{z}\geq 2+d_{z}$, which simplifies to $d_{w}\geq 1$. For the supply equation, we have $2+d_{w}+d_{z}$ instruments and $3+d_{w}$ parameters, necessitating $2+d_{w}+d_{z}\geq 3+d_{q}$, which reduces to $d_{z}\geq 1$.
Under the rank conditions (B.17) and (B.19), when the system is exactly identified, the parameter vectors have unique solutions given by:
$\boldsymbol{\alpha}$ $=\left(\mathbb{E}[\mathbf{Z}_{D,mt}\mathbf{X}_{D,mt}^{\top}]\right)^{-1}\mathbb{E}[\mathbf{Z}_{D,mt}Q_{mt}]$ (B.21)
$\boldsymbol{\theta}$ $=\left(\mathbb{E}[\mathbf{Z}_{S,mt}\mathbf{X}_{S,mt}^{\top}]\right)^{-1}\mathbb{E}[\mathbf{Z}_{S,mt}P_{mt}]$ (B.22)
Proof of Lemma 2 -1.
Two parameter sets $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})$ and $(\theta_{1}^{*},\lambda_{\text{pre}}^{*},\lambda_{\text{post}}^{*})$ are observationally equivalent if they generate the same values for the observables:
$\tilde{\gamma}_{0}$ $=\theta_{1}-\frac{\lambda_{\text{pre}}}{\alpha_{1}}=\theta_{1}^{*}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}}=\tilde{\gamma}_{0}^{*},$ (B.23)
$\tilde{\gamma}_{1}$ $=\theta_{1}-\frac{\lambda_{\text{post}}}{\alpha_{1}}=\theta_{1}^{*}-\frac{\lambda_{\text{post}}^{*}}{\alpha_{1}}=\tilde{\gamma}_{1}^{*}.$ (B.24)
We will prove non-identification by constructing an infinite family of observationally equivalent parameter sets. Let $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})$ be any parameter set that satisfies the above two equations. For any arbitrary constant $\kappa\in\mathbb{R}$, define the alternative parameter set $(\theta_{\text{post}}^{*}=\theta_{\text{post}}+\frac{c}{\alpha_{1}};\lambda_{\text{pre}}^{*}=\lambda_{\text{pre}}+\kappa;\lambda_{\text{post}}^{*}=\lambda_{\text{post}}+c)$. We can verify that this alternative parameter set generates the same observables, and hence, they are observationally equivalent. However, for any $\kappa\neq 0$, we have $\theta_{1}^{*}\neq\theta_{1}$, $\lambda_{\text{pre}}^{*}\neq\lambda_{\text{pre}}$, and $\lambda_{\text{post}}^{*}\neq\lambda_{\text{post}}$. Since $\kappa$ was arbitrarily, the parameters $\theta_{1}$, $\lambda_{\text{pre}}$, and $\lambda_{\text{post}}$ are not identified.
Next, we show that the change in conduct $\Delta\lambda:=\lambda_{\text{post}}-\lambda_{\text{pre}}$ is identified. To this end, we show that $\Delta\lambda$ takes the same value across all observationally equivalent parameter sets. Consider any two observationally equivalent parameter sets $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})$ and $(\theta_{1}^{*},\lambda_{\text{pre}}^{*},\lambda_{\text{post}}^{*})$ that generate the same observables $\tilde{\gamma}_{0}$ and $\tilde{\gamma}_{1}$. It is easy to check that for both these sets, $\Delta\lambda=\lambda_{\text{post}}-\lambda_{\text{pre}}=-\alpha_{1}\times(\tilde{\gamma}_{1}-\tilde{\gamma}_{0})=-\alpha_{1}\times(\tilde{\gamma}_{1}^{*}-\tilde{\gamma}_{0}^{*})=\lambda_{\text{post}}^{*}-\lambda_{\text{pre}}^{*}=\Delta\lambda.$ Therefore, $\Delta\lambda$ takes the same value across all observationally equivalent parameter sets, which proves that $\Delta\lambda$ is identified. ∎
Proof of Lemma 2-2.
Two parameter sets $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})$ and $(\theta_{1}^{*},\lambda_{\text{pre}}^{*},\lambda_{\text{post}}^{*})$ are observationally equivalent if $(\tilde{\gamma}_{0}^{1},\tilde{\gamma}_{1}^{1},\tilde{\gamma}_{0}^{2},\tilde{\gamma}_{1}^{2})=(\tilde{\gamma}_{0}^{1*},\tilde{\gamma}_{1}^{1*},\tilde{\gamma}_{0}^{2*},\tilde{\gamma}_{1}^{2*})$. The parameter vector $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})$ is identified if for any other parameter vector $(\theta_{1}^{*},\lambda_{\text{pre}}^{*},\lambda_{\text{post}}^{*})$ that is observationally equivalent, we have $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})=(\theta_{1}^{*},\lambda_{\text{pre}}^{*},\lambda_{\text{post}}^{*})$. Suppose there are two parameter sets $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})$ and $(\theta_{1}^{*},\lambda_{\text{pre}}^{*},\lambda_{\text{post}}^{*})$ that are observationally equivalent.
So, the following must be true
$\tilde{\gamma}_{0}^{1}=\theta_{1}-\frac{\lambda_{\text{pre}}}{\alpha_{1}^{1}}$ $=\theta_{1}^{*}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{1}}=\tilde{\gamma}_{0}^{1*},$ (B.25)
$\tilde{\gamma}_{1}^{1}=\theta_{1}-\frac{\lambda_{\text{post}}}{\alpha_{1}^{1}}$ $=\theta_{1}^{*}-\frac{\lambda_{\text{post}}^{*}}{\alpha_{1}^{1}}=\tilde{\gamma}_{1}^{1*},$ (B.26)
$\tilde{\gamma}_{0}^{2}=\theta_{1}-\frac{\lambda_{\text{pre}}}{\alpha_{1}^{2}}$ $=\theta_{1}^{*}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{2}}=\tilde{\gamma}_{0}^{2*},$ (B.27)
$\tilde{\gamma}_{1}^{2}=\theta_{1}-\frac{\lambda_{\text{post}}}{\alpha_{1}^{2}}$ $=\theta_{1}^{*}-\frac{\lambda_{\text{post}}^{*}}{\alpha_{1}^{2}}=\tilde{\gamma}_{1}^{2*}.$ (B.28)
Rearranging Equations (B.25) and (B.27) gives $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{1}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{1}}$ and $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{2}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{2}}$, respectively. Equating the two and simplifying the expressions further gives $\lambda_{\text{pre}}\times\left(\frac{\alpha_{1}^{2}-\alpha_{1}^{1}}{\alpha_{1}^{1}\alpha_{1}^{2}}\right)=\lambda_{\text{pre}}^{*}\times\left(\frac{\alpha_{1}^{2}-\alpha_{1}^{1}}{\alpha_{1}^{1}\alpha_{1}^{2}}\right).$ As $\alpha_{1}^{1}\neq\alpha_{1}^{2}$, the two terms in the parentheses in the equation above are non-zero, and therefore $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$. Applying the same logic as above to Equations (B.26) and (B.28), it follows that we get $\lambda_{\text{post}}=\lambda_{\text{post}}^{*}$.
Finally, we note that because $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$ and $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{1}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{1}}=0$, as shown above, we get $\theta_{1}=\theta_{1}^{*}$. Now that we have established $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$ and $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$. This proves that the parameters $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{pre}})$ are identified, and are given by $\theta_{1}=\frac{\alpha_{1}^{1}\tilde{\gamma}_{0}^{1}-\alpha_{1}^{2}\tilde{\gamma}_{0}^{2}}{\alpha_{1}^{1}-\alpha_{1}^{2}};\quad\!\!\lambda_{\text{pre}}=\min\{0,\max\{\alpha_{1}^{1}(\theta_{1}-\tilde{\gamma}_{0}^{1}),1\}\};\quad\!\!\lambda_{\text{post}}=\min\{0,\max\{\alpha_{1}^{1}(\theta_{1}-\tilde{\gamma}_{1}^{1}),1\}\}.$
C Robustness: Fixed Effects Estimates
This appendix examines the robustness of our main findings by exploring fixed effects specifications that we argue in Sections 3.3 and 4.3.2. We demonstrate that the apparently pro-competitive merger effects found using standard DiD with fixed effects disappear when we apply synthetic control methods to create more appropriate comparisons. These results support our preferred first-difference approach and provide more reliable identification.
C.1 Synthetic DiD
Our data (see Table 1) suggests that standard DiD estimates using fixed effects may be misleading, not because treated markets experience larger price declines (average prices went from \$228.25 to \$226.33 post merger), but because control markets experience systematic price increases (from \$185.65 to \$195.75). This pattern suggests that our control markets may not be the right ones.
To address this concern, we implement synthetic difference-in-differences (SDID) following Arkhangelsky et al. (2021), which constructs market- and time-specific weights to create better counterfactuals by matching pre-treatment price trends between treated and control markets. Table C.1 presents the results for both level and log specifications.
Table C.1: Synthetic DiD Treatment Effects
Panel A: Level Panel B: Log
DL–NW UA–CO AA–US DL–NW UA–CO AA–US
Price -0.0232 0.0245 -0.0827 -0.0148 -0.0044 -0.0355
(0.0522) (0.0204) (0.0756) (0.0221) (0.0120) (0.0318)
Passengers 0.0378 -0.1362 0.2615 0.0485 -0.0110 0.0782
(0.0459) (0.1913) (0.2127) (0.0345) (0.0257) (0.0523)
Seats 1.4855∗∗ 1.7849∗∗∗ 1.6218∗∗∗ 0.1296 0.2134∗∗∗ 0.1230∗∗∗
(0.6214) (0.2235) (0.3975) (0.1000) (0.0153) (0.0296)
  • Note: Table reports synthetic DiD treatment effects for three airline mergers using both level and log specifications. Level specification presents results with prices (in ’00 \$) and passenger volumes/seats (in 10,000). Log specification shows log-transformed outcomes. Standard errors in parentheses. Significance levels: ∗∗∗ $p<0.01$, ∗∗ $p<0.05$, $p<0.1$.
The synthetic DiD results, like the first-difference method, contrast with standard fixed effects DiD estimates. Once we construct appropriate synthetic controls that match pre-treatment price trends, the large pro-competitive price effects found in standard DiD with fixed effect analyses vanish entirely. None of the price effects are statistically significant, and most point estimates are economically small. The only consistently significant effects appear for seat capacity, which increases following mergers. These results support our main text argument that standard DiD approaches to merger evaluation may systematically overstate pro-competitive effects.
C.2 Structural Estimates with Fixed Effects
We next examine whether similar patterns emerge in our structural estimation when using fixed effects rather than first differences. Table C.2 compares standard GMM estimates with our synthetic GMM approach, both using a fixed effects specification.
The structural estimates using fixed effects without synthetic weights tell a dramatically different story from our preferred first-difference results. Standard GMM with fixed effects suggests that all three mergers generated significant pro-competitive effects: DL-NW and AA-US show large negative aggregate price effects (-\$34.2 and -\$41.9 per hundred dollars, respectively). At the same time, UA-CO demonstrates positive but significant effects (\$29.0). These results suggest that airline mergers uniformly benefited consumers through substantial price reductions.
Table C.2: Structural Estimation with Fixed Effects
GMM Synthetic GMM
DL-NW UA-CO AA-US DL-NW UA-CO AA-US
Panel A:
Change in Conduct 0.138*** -0.073** 0.146** 0.013 -0.102** 0.195*
(0.036) (0.028) (0.061) (0.030) (0.052) (0.106)
Panel B:
Average Efficiency -0.286*** 0.411** -0.557*** 0.072 0.173* -1.451
(0.056) (0.165) (0.209) (0.182) (0.092) (1.971)
Panel C:
Aggregate Price Effect of Merger -0.342*** 0.290*** -0.419*** 0.034 0.156* -0.215*
(0.064) (0.104) (0.125) (0.081) (0.084) (0.127)
Observations 13,304 12,299 12,458 13,304 12,299 12,458
  • Note: Table presents GMM and synthetic GMM estimates using fixed effects. Panel A presents changes in strategic conduct parameter averaged over eight post-merger quarters. Panel B shows average merger-induced cost efficiency over the same period, measured as cost change per hundred dollars. Panel C reports aggregate price effects of merger. Standard errors in parentheses. $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
However, this finding changes completely when we use a synthetic GMM. Under synthetic GMM, the large, significant effects largely disappear, and the estimates become remarkably similar to our preferred first-difference results from the main text. Most coefficients lose statistical significance, and the magnitude of effects decreases substantially. For instance, the DL-NW aggregate price effect shrinks from -34.2 cents (highly significant) to 3.4 cents (insignificant), while the AA-US effect moderates from -41.9 cents to -21.5 cents.
This pattern demonstrates two crucial insights. First, it confirms that fixed effects estimation systematically biases results toward finding pro-competitive merger effects, consistent with our theoretical concerns about DiD identification in merger contexts. Second, it validates the effectiveness of synthetic control weights in addressing these biases—the synthetic GMM estimates align closely with our first-difference results, providing reassurance that both methods are detecting genuine treatment effects rather than methodological artifacts.